# Integrated Optimal Control for Fast Charging and Active Thermal Management of Lithium-Ion Batteries in Extreme Ambient Temperatures

### v1

@@ III-B State-Feedback Model Predictive Control
Depending on the specific configuration of the battery cooling and heating system and their actuators, the definition of $P_{\mathrm{act}}$’s smoothness could be different. This paper adopts the general form to define smoothness since modeling the actuators is not within this paper’s scope.
@@
The state-feedback MPC algorithm is summarized in Algorithm 1. As shown by Line 2, the system state is observed with a smaller sampling time $\Delta_{\mathrm{s}}$, i.e. $\Delta_{\mathrm{s}}<\Delta_{\mathrm{p}}$. The MPC is computed for every time interval $\Delta_{\mathrm{p}}$ given the present state $\boldsymbol{x}_{k}$ at time $t_{k}$, as shown by Line 3. Then for every time instance $t_{i}$, the input $\boldsymbol{u}_{i}$ is calculated by interpolating the optimal control trajectory from the MPC with zero-order hold, as indicated in Line 8.
@@
2       observe $\boldsymbol{x}_{i}$ and $T_{\mathrm{amb},i}$ at time $t_{i}$ if $t_{i}\ \%\ \Delta_{\mathrm{p}}==0$ then // for every time interval $\Delta_{\mathrm{p}}$
@@ 6      $t_{i}\leftarrow t_{i}+\Delta_{\mathrm{s}}$, $i\leftarrow i+1$
III-C System Parameters
The necessary parameters for the battery dynamics (10) and the constraints (11) - (16) are given by Table I. The system parameters are given in [12], using a 3 Ah Panasonic NCR-18650B LiB battery. The constraint parameters are derived from [28]. $\overline{P}_{\mathrm{heat}}$ and $\overline{P}_{\mathrm{cool}}$ are determined by scaling the power limits in [31]. The mapping $h(\cdot)$ in (1f) is parameterized by a fifth-order polynomial [12], i.e.
@@ III-C System Parameters
The model parameters are validated for charging current up to 4.5 A, whose maximum charging C-rate is 1.5 C. This paper defines the upper bound for charging current as $\overline{I}=3$ A. Note that the unit for all the temperature variables or bounds is Kelvin. But for the sake of readability, the rest of this paper adopts degree Celsius to express the same temperature in Kelvin.
@@
$C_{\mathrm{b}}$ 10037 F $C_{\mathrm{core}}$ 40 F $\underline{\mathrm{SoC}}$ 0%
$C_{\mathrm{s}}$ 973 F $C_{\mathrm{surf}}$ 10 F $\overline{\mathrm{SoC}}$ 100%
$R_{\mathrm{b}}$ 0.019 $\Omega$ $R_{\mathrm{core}}$ 4 $\Omega$ $\underline{I}$ 0 A
$\gamma_{1}$ 0.026 $R_{\mathrm{surf}}$ 7 $\Omega$ $\overline{I}$ 3 A
$\gamma_{2}$ 0.061 $\kappa_{1}$ 30 $\underline{V}$ 0 V
$\gamma_{3}$ 14.36 $\kappa_{2}$ 70 $\overline{V}$ 4.2 V
$\alpha_{0}$ 3.2 $T_{\mathrm{ref}}$ 25 ${}^{\circ}$ $\underline{T}_{\mathrm{core}}$10 ${}^{\circ}$
$\alpha_{1}$ 2.59 $\eta_{\mathrm{act}}$ 87% $\overline{T}_{\mathrm{core}}$ 55 ${}^{\circ}$
$\alpha_{2}$9.003 $\beta_{1}$0.04 $\underline{V}_{\mathrm{b}}$ 0 V
$\alpha_{3}$ 18.87 $\beta_{2}$ 0.08 $\overline{V}_{\mathrm{b}}$ 0.95 V
$\alpha_{4}$17.82 $\underline{P}_{\mathrm{act}}$8 W $\underline{V}_{\mathrm{s}}$ 0 V
$\alpha_{5}$ 6.325 $\overline{P}_{\mathrm{act}}$ 8 W $\overline{V}_{\mathrm{s}}$ 0.95 V
@@ IV Numerical Experiments
This section presents several case studies with numerical experiments to illustrate the effectiveness of the proposed thermal-NDC model in integrated charging and active thermal management under extreme ambient temperatures. Additionally, this section investigates the impact of MPC parameters, battery core temperature, and active thermal power on fast charging. Several insights are revealed and summarized in this section to explain why the decisions made by the proposed algorithm lead to energy-efficient fast charging.
@@ IV-A Basic Case Study
P
the proposed MPC (2), optimization initial guess given by forward propagating dynamics with zero inputs
P1
same as P, optimization initial guess given by forward propagating dynamics with maximum charging current and $P_{\mathrm{act},k}$; $P_{\mathrm{act},k}$ is forwardly calculated by (20) with $T_{\mathrm{core,r}}=45$ ${}^{\circ}\text{C}$
A
the proposed MPC (2) without active thermal management, i.e. $\underline{P}_{\mathrm{act}}=\overline{P}_{\mathrm{act}}=0$ W, $P_{\mathrm{act}}(t)=0,\ \forall t$
B
A, battery temperature is separately controlled by the PID controller (20), $T_{\mathrm{core,r}}=25$ ${}^{\circ}\text{C}$
C
same as B, $T_{\mathrm{core,r}}=35$ ${}^{\circ}\text{C}$
D
same as B, $T_{\mathrm{core,r}}=45$ ${}^{\circ}\text{C}$
E
same as B, $T_{\mathrm{core,r}}=50$ ${}^{\circ}\text{C}$
@@ IV-A1 Mild Ambient Temperature
The ambient temperature $T_{\mathrm{amb}}=25$ ${}^{\circ}\text{C}$. The initial state $\boldsymbol{x}_{0}=\begin{bmatrix}0.1\text{ V}&0.1\text{ V}&25\,^{\circ}\text{C}&25\,^{\circ}\text{C}\end{bmatrix}^{\top}$. The simulation results are shown in Table III, where $T_{\mathrm{chg}}$ indicates the total charging time; $T_{\mathrm{comp}}$ indicates the average computational time and its standard deviation; Energy indicates the total energy used for charging and thermal management; Efficiency indicates the charging efficiency (the ratio of the energy used for raising SoC to the total energy), i.e.
@@
According to Table III, strategies P and P1 have the smallest charging time. As for total consumed energy and efficiency, A outperforms the others since charging the battery without active thermal management will not violate the battery temperature constraint when the ambient temperature is mild. Fig. 2a illustrates the trajectories of battery core and surface temperature and active thermal power for strategies P and A, where strategy P chooses to warm the battery after about 1800 s and thus results in faster charging. From strategies B to D, as the desired battery core temperature $T_{\mathrm{core,r}}$ increasing from 25 ${}^{\circ}\text{C}$ to 45 ${}^{\circ}\text{C}$, the charging time decreases from 3019 s to 3009 s. But when the desired core temperature $T_{\mathrm{core,r}}=50$ ${}^{\circ}\text{C}$ is closed to the upper bound $\overline{T}_{\mathrm{core}}=55$$^{\circ}$, the PID controller violates $\overline{T}_{\mathrm{core}}$ at 690 s - 1090 s since it has no predictive information on how the system dynamics will evolve. These observations are consistent with the statement that heating the battery could benefit the charging speed [6, 7].
@@ IV-A3 Low Ambient Temperature
${}^{\dagger}$ N/A N/A N/A $106.88\pm 63.50$
${}^{\ddagger}$ 3022 57.60 56.24% $21.93\pm 9.20$
${}^{\ddagger\mathsection}$ 3022 57.60 56.24% $21.96\pm 9.05$
${}^{\ddagger\mathsection}$ 3023 47.63 68.01% $24.67\pm 4.85$
${}^{\ddagger\mathsection}$ 3022 57.60 56.24% $21.98\pm 9.07$
@@ IV-A4 Summary on Basic Case Study
The advantage of the proposed thermal-NDC model can be also reflected by an observation from Fig. 2, where $P_{\mathrm{act}}$ drops toward cooling to reduce $T_{\mathrm{core}}$ and $T_{\mathrm{surf}}$ at the end of charging. In fact, as Fig. 3 illustrated, the charging current also drops at the same time. The dropping on both $P_{\mathrm{act}}$ and $I$ is to reduce $V_{\mathrm{s}}$ and $V$ such that they will not violate the constraints as they approach the upper bounds closely. Also, as SOC increases, the upper bound on the Li-ion concentration gradient within the electrode, i.e. $\beta_{1}\mathrm{SoC}(t)+\beta_{2}$, decreases. This again requires $V_{\mathrm{s}}$ to reduce such that $V_{\mathrm{s}}-V_{\mathrm{b}}$ will not violate its upper bound. The system behaviors with $T_{\mathrm{amb}}=-25$ ${}^{\circ}\text{C}$ and $T_{\mathrm{amb}}=25$ ${}^{\circ}\text{C}$ are similar to Fig. 3, thereby omitted.
@@
As for the computational time between P and P1, Tukey’s HSD (honestly significant difference) Test [34, Chapter 14] is performed for three cases and it reveals a statistically significant difference between the computational time of strategies P and P1; the computational time of P is statistically greater than P1’s with the p-value 0.00 $<$ 0.05. This indicates that the initial guess of P1’s MPC problem, i.e. regulating $T_{\mathrm{core}}$ to $T_{\mathrm{core,r}}=45$ ${}^{\circ}\text{C}$ in all three ambient temperatures, is closer to the optimum than P’s initial guess, which benefits the run-time computation.
A similar conclusion can be drawn based on the observation from Fig. 2, where the MPC strategy P seems to maintain a high $T_{\mathrm{core}}$ for fast charging. These two observations are consistent with the conclusions made by [6, 7]. However, compared with the results in mild and high ambient temperatures, it is uncertain that raising the heating power bound could further speed up the charging given low ambient temperature. Also, comparing the computational time of P and P1 in the low and high ambient temperatures, the initialization of P1, i.e. regulating $T_{\mathrm{core}}$ to $T_{\mathrm{core,r}}=45$ ${}^{\circ}\text{C}$, might be closer to the optimal solution than strategy P’s initialization. Thus, the next subsection investigates how battery core temperature and active heating would affect the charging speed, and verify whether an MPC strategy with a long enough horizon can obtain an optimal solution that implicitly encourages warming up the battery to an optimal temperature.
@@ IV-B Other Factors
Table VI: Strategy Description for Other Factors
Name Description
P2
same as ${}^{\dagger}$, objective function using $J_{2}$ in (22), $T_{\mathrm{core,r}}=45^{\circ}\text{C}$
P3
same as P2, $N=40$, $T_{\mathrm{core,r}}=55$ ${}^{\circ}\text{C}$
P4
same as P, $N=80$
P5
same as P, $N=120$
  • $\dagger$
    Note: for strategy P, $N=40$
Regarding the computational time of the proposed strategies P and P1 - P5, a one-way analysis of variance (one-way ANOVA) [34, Chapter 14] is performed, which reveals a statistically significant difference in computational time between at least two groups with a p-value of $0.00<0.05$. Then Tukey’s HSD Test reveals a relation with statistical significance among these groups, i.e. P1 $<$ P $<$ P2 $<$ P3 $<$ P4 $<$ P5.
@@
As for the effect of $T_{\mathrm{core}}$ on charging speed, P3 outperforms the other strategies in terms of charging time. Meanwhile, the other proposed strategies yield slightly longer charging times. Nevertheless, P1 can achieve a good balance across all indices. Fig. 4b illustrates the reason why P3 outperforms the others on charging time, where P3 tracks $T_{\mathrm{core,r}}=55$ ${}^{\circ}\text{C}$ to maintain a high $T_{\mathrm{core}}$ for a longer duration than P1 and P5. The reason can be revealed by the proposed thermal-NDC model, where a higher $T_{\mathrm{core}}$ reduces the internal resistance $R_{\mathrm{b,T}}(t)$ as (5) describes, and further affects the capacitor voltage dynamics (1e). Since $\mathrm{SoC}(t)$ is defined as a linear combination of the capacitor voltages $V_{\mathrm{b}}(t)$ and $V_{\mathrm{s}}(t)$ in (3), the time derivative of SOC is directly affected by the capacitor voltage dynamics (1e). This explanation is consistent with the observation from numerical experiments, as well as the conclusions from the literature [6, 7]. However, heating the battery at the beginning as P3 does is not energy-efficient. Fig. 4a and Fig. 4c indicate that after the battery is charged for about 2200 s and 1700 s, P1 and P5 begin to heat the battery because they generate more heat by (7) with a higher $V$.

### v3

@@ III-B State-Feedback Model Predictive Control
The required smoothness of the charging current imposed by (18) may vary or even be unnecessary, depending on the specifications of the power electronics. For tutorial purposes, this paper adopts a general form to define the smoothness requirement as part of the objective, but it can be adjusted to meet specific needs. The smoothness requirement for $P_{\mathrm{act}}$ in (18) is imposed for similar reasons.
@@
The state-feedback MPC algorithm is summarized in Algorithm 1. As shown by Line 2, the system state is observed with a smaller sampling time $\Delta_{\mathrm{s}}$, i.e., $\Delta_{\mathrm{s}}<\Delta_{\mathrm{p}}$. The MPC is computed for every time interval $\Delta_{\mathrm{p}}$ given the present state $\boldsymbol{x}_{k}$ at time $t_{k}$, as shown by Line 3. Then for every time instance $t_{i}$, the input $\boldsymbol{u}_{i}$ is calculated by interpolating the optimal control trajectory from the MPC with zero-order hold, as indicated in Line 8.
@@
2       observe $\boldsymbol{x}_{i}$ and $T_{\mathrm{amb},i}$ at time $t_{i}$
3
4      if $t_{i}\ \%\ \Delta_{\mathrm{p}}==0$ then // for every time interval $\Delta_{\mathrm{p}}$
5             $k\leftarrow k+1$
@@ 11      $t_{i}\leftarrow t_{i}+\Delta_{\mathrm{s}}$, $i\leftarrow i+1$
III-C System Parameters and Identification
The necessary parameters for the battery dynamics (10) and the constraints (11) - (16) are given by Table I. The system parameters are given in [12], using a 3 Ah Panasonic NCR-18650B LiB battery. The constraint parameters are derived from [28]. $\overline{P}_{\mathrm{heat}}$ and $\overline{P}_{\mathrm{cool}}$ are determined by scaling the power limits in [44]. The mapping $h(\cdot)$ in (1b) is parameterized by a fifth-order polynomial [12], i.e.
@@ III-C System Parameters and Identification
The model parameters are validated for charging current up to 4.5 A, whose maximum charging C-rate is 1.5C. This paper defines the upper bound for charging current as $\overline{I}=3$ A. For the sake of readability, the rest of this paper adopts degree Celsius to express the same temperature in Kelvin.
To identify and calibrate the battery parameters, a two-step procedure is adopted as described in [12]. This process involves initially identifying the $h(\cdot)$ mapping and subsequently estimating the impedance and capacitance parameters.
First, the SoC-OCV relationship of the proposed model is given by $\text{OCV}=h(\text{SoC})$. The $h(\cdot)$ mapping can be identified by fitting it with a battery’s SoC-OCV data. To obtain this data, the battery is discharged at a low current (e.g., 1/25 C-rate) from full to empty. During this process, the terminal voltage $V$ is taken as OCV, and the SoC is calculated using the Coulomb counting method [12]. The mapping $h(\cdot)$ is then identified as a parameterized polynomial by solving a least squares problem.
Second, to identify the impedance and capacitance parameters, the battery is discharged at a constant current of normal magnitude. These parameters are estimated by expressing the terminal voltage in terms of the parameters and fitting this expression to the measurement data [12]. For further details, refer to [12, 45].
@@
$C_{\mathrm{b}}$ 10037 F $C_{\mathrm{core}}$ 40 J/K $\underline{\mathrm{SoC}}$ 0%
$C_{\mathrm{s}}$ 973 F $C_{\mathrm{surf}}$ 10 J/K $\overline{\mathrm{SoC}}$ 100%
$R_{\mathrm{b}}$ 0.019 $\Omega$ $R_{\mathrm{core}}$ 4 K/W $\underline{I}$ 0 A
$\gamma_{1}$ 0.026 $\Omega$ $R_{\mathrm{surf}}$ 7 K/W $\overline{I}$ 3 A
$\gamma_{2}$ 0.061 $\Omega$ $\kappa_{1}$ 30 K $\underline{V}$ 0 V
$\gamma_{3}$ 14.36 $\kappa_{2}$ 70 K $\overline{V}$ 4.2 V
$\alpha_{0}$ 3.2 V $T_{\mathrm{ref}}$ 298.15 K $\underline{T}_{\mathrm{core}}$ 263.15 K
$\alpha_{1}$ 2.59 $\eta_{\mathrm{act}}$ 87% $\overline{T}_{\mathrm{core}}$ 328.15 K
$\alpha_{2}$9.003 V-1 $\beta_{1}$0.04 V $\underline{V}_{\mathrm{b}}$ 0 V
$\alpha_{3}$ 18.87 V-2 $\beta_{2}$ 0.08 V $\overline{V}_{\mathrm{b}}$ 1 V
$\alpha_{4}$17.82 V-3 $\underline{P}_{\mathrm{act}}$8 W $\underline{V}_{\mathrm{s}}$ 0 V
$\alpha_{5}$ 6.325 V-4 $\overline{P}_{\mathrm{act}}$ 8 W $\overline{V}_{\mathrm{s}}$ 1 V
@@ IV Numerical Experiments
This section presents several case studies with numerical experiments to illustrate the effectiveness of the proposed thermal-NDC model in integrated charging and active thermal management under extreme ambient temperatures. Additionally, this section investigates the impact of MPC parameters, battery core temperature, and active thermal power on fast charging. Several insights are revealed and summarized in this section to explain why the decisions made by the proposed algorithm lead to energy-efficient and health-aware optimal fast charging.
@@ IV-A Basic Case Study
P the proposed MPC (4), optimization initial guess given by forward propagating dynamics with zero inputs
P1 same as P, optimization initial guess given by forward propagating dynamics with maximum charging current and $P_{\mathrm{act},k}$; $P_{\mathrm{act},k}$ is forwardly calculated by (20) with $T_{\mathrm{core,r}}=45$ ${}^{\circ}\text{C}$
A the proposed MPC (4) without active thermal management, i.e., $\underline{P}_{\mathrm{act}}=\overline{P}_{\mathrm{act}}=0$ W, $P_{\mathrm{act}}(t)=0,\ \forall t$
B A, battery temperature is separately controlled by the PID controller (20), $T_{\mathrm{core,r}}=25$ ${}^{\circ}\text{C}$
C same as B, $T_{\mathrm{core,r}}=35$ ${}^{\circ}\text{C}$
D same as B, $T_{\mathrm{core,r}}=45$ ${}^{\circ}\text{C}$
E same as B, $T_{\mathrm{core,r}}=50$ ${}^{\circ}\text{C}$
@@ IV-A1 Mild Ambient Temperature
The ambient temperature $T_{\mathrm{amb}}=25$ ${}^{\circ}\text{C}$. The initial state $\boldsymbol{x}_{0}=\begin{bmatrix}0.1\text{ V}&0.1\text{ V}&25\,^{\circ}\text{C}&25\,^{\circ}\text{C}\end{bmatrix}^{\top}$. The simulation results are shown in Table III, where $T_{\mathrm{chg}}$ indicates the total charging time; $T_{\mathrm{comp}}$ indicates the average computational time and its standard deviation; Energy indicates the total energy used for charging and thermal management; Efficiency indicates the charging efficiency (the ratio of the energy used for raising SoC to the total energy), i.e.,
@@
According to Table III, strategies P and P1 have the smallest charging time. As for total consumed energy and efficiency, A outperforms the others since charging the battery without active thermal management will not violate the battery temperature constraint when the ambient temperature is mild. Fig. 2a illustrates the trajectories of battery core and surface temperature and active thermal power for strategies P and A, where strategy P chooses to warm the battery after about 1800 s and thus results in faster charging. From strategies B to D, as the desired battery core temperature $T_{\mathrm{core,r}}$ increasing from 25 ${}^{\circ}\text{C}$ to 45 ${}^{\circ}\text{C}$, the charging time decreases from 3019 s to 3009 s. But when the desired core temperature $T_{\mathrm{core,r}}=50$ ${}^{\circ}\text{C}$ is closed to the upper bound $\overline{T}_{\mathrm{core}}=55$ C, the PID controller violates $\overline{T}_{\mathrm{core}}$ at 690 s - 1090 s since it has no predictive information on how the system dynamics will evolve. These observations are consistent with the statement that heating the battery could benefit the charging speed [6, 7].
@@ IV-A3 Low Ambient Temperature
A N/A N/A N/A $106.88\pm 63.50$
B 3022 57.60 56.24% $21.93\pm 9.20$
C‡§ 3022 57.60 56.24% $21.96\pm 9.05$
D‡§ 3023 47.63 68.01% $24.67\pm 4.85$
E‡§ 3022 57.60 56.24% $21.98\pm 9.07$
@@ IV-A4 Summary on Basic Case Study
The advantage of the proposed thermal-NDC model can be also reflected by an observation from Fig. 2, where $P_{\mathrm{act}}$ drops toward cooling to reduce $T_{\mathrm{core}}$ and $T_{\mathrm{surf}}$ at the end of charging. In fact, as Fig. 3 illustrated, the charging current also drops at the same time. The dropping on both $P_{\mathrm{act}}$ and $I$ is to reduce $V_{\mathrm{s}}$ and $V$ such that they will not violate the constraints as they approach the upper bounds closely. Also, as SoC increases, the upper bound on the Li-ion concentration gradient within the electrode, i.e., $\beta_{1}\mathrm{SoC}(t)+\beta_{2}$, decreases. This again requires $V_{\mathrm{s}}$ to reduce such that $V_{\mathrm{s}}-V_{\mathrm{b}}$ will not violate its upper bound. The system behaviors with $T_{\mathrm{amb}}=-25$ ${}^{\circ}\text{C}$ and $T_{\mathrm{amb}}=25$ ${}^{\circ}\text{C}$ are similar to Fig. 3, thereby omitted.
@@
As for the computational time between P and P1, Tukey’s HSD (honestly significant difference) Test [48, Chapter 14] is performed for three cases and it reveals a statistically significant difference between the computational time of strategies P and P1; the computational time of P is statistically greater than P1’s with the p-value 0.00 $<$ 0.05. This indicates that the initial guess of P1’s MPC problem, i.e., regulating $T_{\mathrm{core}}$ to $T_{\mathrm{core,r}}=45$ ${}^{\circ}\text{C}$ in all three ambient temperatures, is closer to the optimum than P’s initial guess, which benefits the run-time computation.
A similar conclusion can be drawn based on the observation from Fig. 2, where the MPC strategy P seems to maintain a high $T_{\mathrm{core}}$ for fast charging. These two observations are consistent with the conclusions made by [6, 7]. However, compared with the results in mild and high ambient temperatures, it is uncertain that raising the heating power bound could further speed up the charging given low ambient temperature. Also, comparing the computational time of P and P1 in the low and high ambient temperatures, the initialization of P1, i.e., regulating $T_{\mathrm{core}}$ to $T_{\mathrm{core,r}}=45$ ${}^{\circ}\text{C}$, might be closer to the optimal solution than strategy P’s initialization. Thus, the next subsection investigates how battery core temperature and active heating would affect the charging speed, and verify whether an MPC strategy with a long enough horizon can obtain an optimal solution that implicitly encourages warming up the battery to an optimal temperature.
@@ IV-B Other Factors
Table VI: Strategy Description for Other Factors
Name Description
P2 same as P, objective function using $J_{2}$ in (22), $T_{\mathrm{core,r}}=45^{\circ}\text{C}$
P3 same as P2, $N=40$, $T_{\mathrm{core,r}}=55$ ${}^{\circ}\text{C}$
P4 same as P, $N=80$
P5 same as P, $N=120$
  • $\dagger$
    Note: for strategy P, $N=40$
@@
Regarding the computational time of the proposed strategies P and P1 - P5, a one-way analysis of variance (one-way ANOVA) [48, Chapter 14] is performed, which reveals a statistically significant difference in computational time between at least two groups with a p-value of $0.00<0.05$. Then Tukey’s HSD Test reveals a relation with statistical significance among these groups, i.e., P1 $<$ P $<$ P2 $<$ P3 $<$ P4 $<$ P5.
@@
As for the effect of $T_{\mathrm{core}}$ on charging speed, P3 outperforms the other strategies in terms of charging time. Meanwhile, the other proposed strategies yield slightly longer charging times. Nevertheless, P1 can achieve a good balance across all indices. Fig. 4b illustrates the reason why P3 outperforms the others on charging time, where P3 tracks $T_{\mathrm{core,r}}=55$ ${}^{\circ}\text{C}$ to maintain a high $T_{\mathrm{core}}$ for a longer duration than P1 and P5. The reason can be revealed by the proposed thermal-NDC model, where a higher $T_{\mathrm{core}}$ reduces the internal resistance $R_{\mathrm{b,T}}(t)$ as (5) describes, and further affects the capacitor voltage dynamics (1a). Since $\mathrm{SoC}(t)$ is defined as a linear combination of the capacitor voltages $V_{\mathrm{b}}(t)$ and $V_{\mathrm{s}}(t)$ in (3), the time derivative of SoC is directly affected by the capacitor voltage dynamics (1a). This explanation is consistent with the observation from numerical experiments, as well as the conclusions from the literature [6, 7]. However, heating the battery at the beginning as P3 does is not energy-efficient. Fig. 4a and Fig. 4c indicate that after the battery is charged for about 2200 s and 1700 s, P1 and P5 begin to heat the battery because they generate more heat by (7) with a higher $V$.